Solve : x^2 - 11x = -28
Answer by stanbon(75887)

Solve : x^2 - 11x = -28
-----
x^2 - 11x + 28 = 0
---
(x-7)(x-4) = 0
x = 7 or x = 4
---------------
Cheers,
Stan H.
----------